OpenStudy (anonymous):

Okay I'm almost finished, Can you guys check this one for me?? Thank YOU!! 12. What is an equation for the line with slope 2/3 and y-intercept 9? a.y=2/3x <-- b. y=9x c.y=2/3x+9 d. y=9x +2/3 I think I have got it wrong but i'm not sure and the other answers don't make sense with the question to me?

OpenStudy (fibonaccichick666):

do you know slope-intercept form?

OpenStudy (cwrw238):

the equation will be of the form y = mx + b where m = slope and b = intercept on the y axis

OpenStudy (anonymous):

It is actually C. y=2/3*x+9 Multiply (2)/(3) by x to get (2x)/(3). y=(2x)/(3)+9 To find the slope and y intercept, use the y=mx+b formula where m=slope and b is the y intercept. y=mx+b Using the y=mx+b formula, m=(2)/(3) and b=9. m=(2)/(3), b=9

OpenStudy (anonymous):

um i learned it but i don't understand it at all

OpenStudy (anonymous):

Thank you all of you!

OpenStudy (anonymous):

welcome :D

OpenStudy (cwrw238):

yw

OpenStudy (fibonaccichick666):

yw, but what part don't you understand? it is a really important concept so maybe we can help you with it

OpenStudy (anonymous):

well i know when you are trying to find the slope on a graph you do rise over run and see how far each point is away from each other. Is that right?

OpenStudy (fibonaccichick666):

yep, so if i gave you two points, could you calculate the slope?

OpenStudy (cwrw238):

the line will look something like this